# DECIPHERING THE DEFINITION OF ADVERSARIAL ROBUSTNESS FOR POST-HOC OOD DETECTORS

### v2

@@ Abstract
Detecting out-of-distribution (OOD) inputs is critical for safely deploying deep learning models in real-world scenarios. In recent years, many OOD detectors have been developed, and even the benchmarking has been standardized, i.e. OpenOOD. The number of post-hoc detectors is growing fast and showing an option to protect a pre-trained classifier against natural distribution shifts, claiming to be ready for real-world scenarios. However, its efficacy in handling adversarial examples has been neglected in the majority of studies. This paper investigates the adversarial robustness of the 16 post-hoc detectors on several evasion attacks and discuss a roadmap towards adversarial defense in OOD detectors.
@@ 1 INTRODUCTION
Current standardized benchmarks such as OpenOOD [5] and RoboDepth [6] merely focus on natural distribution shifts and corruptions [7]. Especially OpenOOD aims to make a fair comparison across methods initially developed for anomaly detection, model uncertainty, open set recognition, and OOD detection methods.
@@
KLM ICML22 $\checkmark$
MLS ICML22 $\checkmark$
@@
Current OOD detection methods, as some listed in Table 1, achieve outstanding results on prominent OOD benchmarks, such as the OpenImage-O [9], ImageNet-O [10], Texture [11], and iNaturalist [12, 13]. OOD detection is a very quickly growing field due to the number of methods added to OpenOOD. More specific post-hoc methods with their plug-and-play capabilities on pre-trained classifiers are more flexible and scalable compared to methods that require full retraining on new OOD data [14, 15]. Simple post-hoc methods like KNN [16] are highlighted maintaining good performance on toy-datasets (e.g. MNIST [17], CIFAR-10 or CIAR-100 [18], and also show outstanding performance on the more realistic dataset like ImageNet [19] according to [14]. These experiments do neglect the adversarial robustness of “state-of-the-art” detectors and the real-world capabilities are questionable as past studies had shown [3, 20, 21, 22]. Adversarial examples remain challenging because they share the same semantics as the training data, but aim to modify the classifier’s output.
@@ 2.1 Evasion Attacks Crafting Inliers
More formally, for an input ${\bm{x}}$ with the ground-truth label $y$, an adversarial example ${\bm{x}}^{\prime}$ is crafted by adding small noise $\bm{\delta}$ to ${\bm{x}}$ such that the predictor model loss ${\bm{J}}({\bm{x}}^{\prime},y)$ is maximized. The $L^{p}$ norm of the adversarial noise should be less than a specified value $\epsilon$, i.e., $\left\|{\bm{x}}-{\bm{x}}^{\prime}\right\|\leq\epsilon$, e.g., $\epsilon=8/255$ [26], to ensure that the image does not change semantically. The attack method Fast Gradient Sign Method (FGSM) by [27] maximizes the loss function in a single step by taking a step towards the sign of the gradient of ${\bm{J}}({\bm{x}},y)$ w.r.t. to ${\bm{x}}$: ${\bm{x}}^{\prime}={\bm{x}}+\epsilon\text{sign}(\nabla_{{\bm{x}}}{\bm{J}}({\bm{x}}^{\prime},y)),$ where the noise meets the $L^{\infty}$ norm bound $\epsilon$. Furthermore, this approach can be applied iteratively, as shown by [28], using a reduced step size $\alpha$: ${\bm{x}}^{\prime}_{0}={\bm{x}},\;{\bm{x}}^{\prime}_{t+1}={\bm{x}}^{\prime}_{t}+\alpha\epsilon\text{sign}(\nabla_{{\bm{x}}}{\bm{J}}({\bm{x}}^{\prime}_{t},y))$, where in each step, the perturbation should be constrained within the $L^{\infty}$ ball of radius $\epsilon$. This constraint is a characteristic of the Projected Gradient Descent (PGD) attack proposed by [1], which is commonly considered a standard attack for the evaluation of model robustness [29]. Masked PGD (mPGD) [30] is a variant of the PGD attack that restricts perturbations to a specific region within an image. The PGD attack is extended towards a patch: ${\bm{x}}_{t+1}^{\prime}=\text{Clip}_{{\bm{x}}^{\prime}}\left({\bm{x}}_{t}^{\prime}+\alpha\cdot\text{sign}(\nabla_{{\bm{x}}}{\bm{J}}({\bm{x}}_{t}^{\prime},y,\bm{\theta})[patch]),\epsilon\right).$ In this context, the term “patch” refers to the region specified as $[x:x+h,y:y+w]$, where $[x,y,h,w]$ are the provided patch’s coordinates and dimensions.
@@ 2.2 Advantages of Post-Hoc OOD Detectors
Post-hoc OOD detection methods, which make use of specific layers of the pre-trained classifier, have been demonstrated to outperform retraining-based approaches, thereby underscoring their empirical efficacy [5]. Their plug-and-play nature allows seamless integration with pre-trained models, without necessitating alterations to the training procedure or access to original training data [5]. In Table 1 are 16 post-hoc detectors listed, which also shows if a method uses features, logits, or the probabilities of the pre-trained model. Post-hoc methods underline their simplicity to outperform others on natural distribution shifts datasets by being as lightweight as possible. The latest post-hoc OOD detector in Table 1 is SCALE [32]. In contrast to the previous activation shaping method, ASH [33], that involves pruning and scaling of activations, SCALE demonstrates state-of-the-art results are solely derived from scaling.
@@
Moreover, post-hoc OOD methods can be augmented with other techniques, such as those employed in OSR [35] or uncertainty estimation [36]. Combining different techniques means also the post-hoc methods become more complex and thus the attack surface might become larger, i.e. attacks against uncertainty estimation [37] differ from the evasion attacks. It is not necessarily the case that a post-hoc method must be combined with other techniques. This demonstrates SAFE [38].
@@ 2.3 OOD Adversarial Detection
3 ROBUSTNESS DEFINITION
@@
The robustness definition in the field of OOD detectors has been ambiguous when it comes to attack methods. There are two categories of adversarial examples. The first one merely attacks the underlying pre-trained classifier and the second one aims to fool the OOD detector itself. Adversarial robustness can be considered in a classifier (Unified Robustness) or the OOD detector (Robust OOD Detection) according to [53]. In this work, we focus on unified robustness, which belong to the covariate shift. For image classification, a dataset $\mathcal{D}=\{({\bm{x}}_{i},y_{i});{\bm{x}}_{i}\in{\mathcal{X}},y_{i}\in{\mathcal{Y}}\}$ sample from a training distribution $\hat{P}_{\rm{data}}({\bm{x}},y)$ is used to train some classifier $C:{\mathcal{X}}\rightarrow{\mathcal{Y}}$. In real-world deployments, distribution shift occurs when classifier $C$ receives data from test distribution $\hat{P}_{\rm{test}}({\bm{x}},y)$ where $\hat{P}_{\rm{data}}({\bm{x}},y)\neq\hat{P}_{\rm{test}}({\bm{x}},y)$ [54]. An OOD detector is a scoring function $s$ that maps an image ${\bm{x}}$ to a real number $\mathbb{R}$ such that some threshold $\tau$ arrives at the detection rule $f({\bm{x}}):\text{ID if }s({\bm{x}})\geq\tau\text{, OOD otherwise}$. Table 4 in Appendix A gives an overview of several detectors as well as considered ID and OOD datasets and model architectures. The ImageNet-1K dataset together with ResNet-50 architecture has become standard for ID. Popular OD datasets are iNaturalist, SUN, Places, and Textures. Some OOD detectors, such as ALOE, OSAD, ADT, ATOM (see Section 2.3), aim to be adversarial robust. They usually take evading attacks (see Section 2.1) such as FGSM or PGD as OOD. These computationally expensive methods only show empirical results on the small-scale CIFAR-10. A robust OOD detector is built to distinguish whether a perturbed input is OOD. Standardized OOD benchmark frameworks, i.e. OpenOOD [5] or RoboDepth [6] do not include unified robustness in their benchmarks at the moment. Consequently, both frameworks, give a false sense of encompassing open-world capabilities. They focus on natural distribution [10], where OOD detection in large-scale semantic space has attracted increasing attention [55], see Appendix A. Some OOD datasets have issues, where ID classes are part of the OD dataset [56]. Recently, [57] found a clean semantic shift dataset that minimizes the interference of covariate shift. The experiments show that state-of-the-art OOD detectors are more sensitive to covariate shift and the advances in semantic shift detection are minimal. The investigation of adversarial examples could gain insights into understanding the covariate shift and towards a generalized OOD detection [4]. The difference between benign ${\bm{x}}$ and the attacked counterpart ${\bm{x}}^{\prime}$ is the different attention of the pre-trained classifier $C$ per sample. Let’s define an attention map ${\bm{A}}$ [58]: The input image is passed through a classifier $C$ to obtain a feature map ${\bm{F}}$. A possible tool could be Grad-CAM [59] to investigate the attention change between benign ${\bm{x}}$ and adversarial example ${\bm{x}}^{\prime}$ on ${\bm{F}}$ [60].
@@ 4 EXPERIMENTS
Experiment Setup
We extend the OpenOOD framework [5] to consider adversarial attacks. We attack the pre-trained classifiers on the corresponding test sets and evaluate 16 post-hoc OOD detectors. As attack methods, we choose FGSM(-$L^{\infty}$), PGD(-$L^{\infty}$), DF(-$L^{2}$) from FoolBox [61]; and the mPGD-($L^{\infty}$). The attacked models are ResNet-18 [62], ResNet-50 [62], and Swin-T [63]. The datasets are CIFAR-10 & CIFAR-100 [18], ImageNet-1K [19] and its variant ImageNet-200 with just 200 classes.
Table 2: Setup. The attack success rate (ASR) from various attacks on different models and datasets. $\mathcal{A}_{\text{std}}$ refers to the standard accuracy of the pre-trained classifier.

### v3

@@ Abstract
Detecting out-of-distribution (OOD) inputs is critical for safely deploying deep learning models in real-world scenarios. In recent years, many OOD detectors have been developed, and even the benchmarking has been standardized, i.e. OpenOOD. The number of post-hoc detectors is growing fast and showing an option to protect a pre-trained classifier against natural distribution shifts, claiming to be ready for real-world scenarios. However, its efficacy in handling adversarial examples has been neglected in the majority of studies. This paper investigates the adversarial robustness of 16 post-hoc detectors on several evasion attacks and discusses a roadmap toward adversarial defense in OOD detectors.
@@ 1 INTRODUCTION
Current standardized benchmarks such as OpenOOD [5] and RoboDepth [6] merely focus on natural distribution shifts and corruptions [7]. Especially OpenOOD aims to make a fair comparison across methods initially developed for anomaly detection [8], model uncertainty [9], open-set recognition (OSR) [10], and OOD detection [11] but neglecting adversarial examples.
@@
KLM ICML’22 $\checkmark$
MLS ICML’22 $\checkmark$
@@
Current OOD detection methods, as some listed in Table 1, achieve outstanding results on prominent OOD benchmarks, such as the OpenImage-O [13], ImageNet-O [14], Texture [15], and iNaturalist [16, 17]. OOD detection is a very quickly growing field by the number of methods added to OpenOOD. More specific post-hoc methods with their plug-and-play capabilities on pre-trained classifiers are more flexible and scalable compared to methods that require full retraining on new OOD data [18, 19]. Simple post-hoc methods like KNN [20] are highlighted maintaining good performance on toy-datasets (e.g. MNIST [21], CIFAR-10 or CIAR-100 [22]), and also show outstanding performance on the more realistic dataset like ImageNet [23] according to [18]. These experiments do neglect the adversarial robustness of “state-of-the-art” (sota) detectors and the real-world capabilities [24] are questionable as past studies had shown [3, 25, 26, 27]. OOD detectors should be able to reliably detect inputs that come from a different distribution than the training data, even if the shift is subtle or complex. Adversarial examples remain challenging because they share the same semantics as the training data, but aim to modify the classifier’s output.
@@ 2 RELATED WORK
This section discusses existing white-box evasion attacks, the benefits of post-hoc detectors, existing methods for detecting adversarial samples, and finally Grad-CAM, a tool for understanding the neural networks’ attention shift initiated by adversarial examples.
@@ 2.1 Evasion Attacks Crafting Inliers
More formally, for an input ${\bm{x}}$ with the ground-truth label $y$, an adversarial example ${\bm{x}}^{\prime}$ is crafted by adding small noise $\bm{\delta}$ to ${\bm{x}}$ such that the predictor model loss ${\bm{J}}({\bm{x}}^{\prime},y)$ is maximized. The $L^{p}$ norm of the adversarial noise should be less than a specified value $\epsilon$, i.e., $\left\|{\bm{x}}-{\bm{x}}^{\prime}\right\|\leq\epsilon$, e.g., $\epsilon=8/255$ [31], to ensure that the image does not change semantically. The attack method Fast Gradient Sign Method (FGSM) by [32] maximizes the loss function in a single step by taking a step towards the sign of the gradient of ${\bm{J}}({\bm{x}},y)$ w.r.t. to ${\bm{x}}$: ${\bm{x}}^{\prime}={\bm{x}}+\epsilon\text{sign}(\nabla_{{\bm{x}}}{\bm{J}}({\bm{x}}^{\prime},y))$, where the noise meets the $L^{\infty}$ norm bound $\epsilon$. Furthermore, this approach can be applied iteratively, as shown by [33], using a reduced step size $\alpha$: $vx^{\prime}_{0}={\bm{x}},\;{\bm{x}}^{\prime}_{t+1}={\bm{x}}^{\prime}_{t}+\alpha\epsilon\text{sign}(\nabla_{{\bm{x}}}{\bm{J}}({\bm{x}}^{\prime}_{t},y))$, where in each step, the perturbation should be constrained within the $L^{\infty}$ ball of radius $\epsilon$. This constraint is a characteristic of the Projected Gradient Descent (PGD) attack proposed by [1], which is commonly considered a standard attack for the evaluation of model robustness [34]. Masked PGD (mPGD) [35] is a variant of the PGD attack that restricts perturbations to a specific region within an image. The PGD attack is extended towards a patch: ${\bm{x}}_{t+1}^{\prime}=\text{Clip}_{{\bm{x}}^{\prime}}\left({\bm{x}}_{t}^{\prime}+\alpha\cdot\text{sign}(\nabla_{{\bm{x}}}{\bm{J}}({\bm{x}}_{t}^{\prime},y,\bm{\theta})[patch]),\epsilon\right).$ In this context, the term “patch” refers to the region specified as $[x:x+h,y:y+w]$, where $[x,y,h,w]$ are the provided patch’s coordinates and dimensions.
@@ 2.2 Advantages of Post-Hoc OOD Detectors
Post-hoc OOD detection methods, which make use of specific layers of the pre-trained classifier, have been demonstrated to outperform retraining-based approaches, thereby underscoring their empirical efficacy [5]. Their plug-and-play nature allows seamless integration with pre-trained models, without necessitating alterations to the training procedure or access to original training data [5]. In Table 1 are 16 post-hoc detectors listed, which also shows if a method uses features, logits, or the probabilities of the pre-trained model. Post-hoc methods underline their simplicity to outperform others on natural distribution shifts datasets by being as lightweight as possible. The latest post-hoc OOD detector in Table 1 is SCALE [37]. In contrast to the previous activation shaping method, ASH [38], that involves pruning and scaling of activations, SCALE demonstrates sota results are solely derived from scaling.
@@
Moreover, post-hoc OOD methods can be augmented with other techniques, such as those employed in OSR [40] or uncertainty estimation [41]. Combining different techniques means also the post-hoc methods become more complex and thus the attack surface might become larger, i.e. attacks against uncertainty estimation [9] differ to the evasion attacks. It is not necessarily the case that a post-hoc method must be combined with other techniques. This demonstrates SAFE [42].
@@ 2.3 OOD Adversarial Detection
2.4 CAM-based Explanations and Consistent Shift by Adversarial Examples
Explainable AI (XAI) [57] can be used as a defense mechanism [58, 59, 60] to spot adversarial examples [61]. The Grad-CAM heatmap provides a clear indication, for a given image, of the regions that the Convolutional Neural Network (CNN) focuses on to make a decision. Class Activation Maps (CAMs) can be considered saliency maps specific to the input [62]. These maps emerge from the accumulated and up-scaled activations at a particular convolutional layer, typically the penultimate layer. The classification is estimated through a linear combination of the activation of units in the final layer(s) of the feature selection network:
$f_{\theta}(.)\approx\sum_{i}\sum_{k}w_{k}a_{ki},$ (1)
where $a_{ki}$ represents the activation of the k-th channel of unit $i$, and $w_{k}$ denotes the learned weights. The relevance values are subsequently defined as $r_{i}=\sum_{k}w_{k}a_{ki}$. The determination of these weights depends on the specific CAM variant employed [63, 64, 65]. Grad-CAM weights the activations using gradients:
$w_{K}:=\frac{\partial f_{\theta}(.)}{\partial a_{ki}}.$ (2)
This weighting directly links to more fundamental explanations that merely estimate the influence of the input on the final output as described before $r_{i}=\partial f_{\theta}(.)/\partial{\bm{x}}_{i}$ [66].
In context to adversarial examples, Chakraborty et al. [67] observes a consistent shift in the region highlighted in the gradient activation heatmap, reflecting its contribution to decision-making across all models under adversarial attacks.
3 DEFINITIONS IN OOD DETECTION
Table 2: Overview of ID and OOD definition from several OOD detectors. The post-hoc detectors are divided into the following categories [5]: Classifiction-based, Density-based, Distance-based. We also mark: Supervised and Adversarial Robust, which are not post-hoc detectors.
Methods ID OOD Model Architectures
Post-Hoc OOD Detection Methods
KNN [20] ImageNet-1K iNaturalist, Places, SUN, Textures ResNet-50
EBO [75] CIFAR-10 ISUN, Places, Texture, SVHN, LSUN WideResNet
ODIN [76] CIFAR-10 LSUN, SVHN, TinyImageNet DenseNet, ResNet
MSP [12] CIFAR-10 SUN (Gaussian) WideResNet 40-4
Adversarial Robust OOD Detectors
ALOE [26] CIFAR-10, CIFAR-100, GSTRB PGD attack DenseNet
OSAD [55] CIFAR-10, SVHN, TinyImageNet FGSM, PGD attack ResNet-18
ADT [46] CIFAR-10, CIFAR-100 FGSM, PGD attack ViT
ATOM [56] CIFAR-10, CIFAR-100, SVHN PGD attack WideResNet
SAFE [42] PASCAL-VOC, DeepDrive FGSM attack RegNetX4.0, ResNet-50
The robustness definition in the field of OOD detectors has been ambiguous when it comes to attack methods. There are two categories of adversarial examples. The first one merely attacks the underlying pre-trained classifier and the second one aims to fool the OOD detector itself. Adversarial robustness can be considered in a classifier (Unified Robustness) or the OOD detector (Robust OOD Detection) according to [77]. This work focuses on unified robustness, which belongs to the covariate shift. For image classification, a dataset $\mathcal{D}=\{({\bm{x}}_{i},y_{i});{\bm{x}}_{i}\in{\mathcal{X}},y_{i}\in{\mathcal{Y}}\}$ sample from a training distribution $\hat{P}_{\rm{data}}({\bm{x}},y)$ is used to train some classifier $C:{\mathcal{X}}\rightarrow{\mathcal{Y}}$. In real-world deployments, distribution shift occurs when classifier $C$ receives data from test distribution $\hat{P}_{\rm{test}}({\bm{x}},y)$ where $\hat{P}_{\rm{data}}({\bm{x}},y)\neq\hat{P}_{\rm{test}}({\bm{x}},y)$ [78]. An OOD detector is a scoring function $s$ that maps an image ${\bm{x}}$ to a real number $\mathbb{R}$ such that some threshold $\tau$ arrives at the detection rule $f({\bm{x}}):\text{ID if }s({\bm{x}})\geq\tau\text{, OOD otherwise}$.
Existing Definitions
@@ 3 DEFINITIONS IN OOD DETECTION
In Table 2, we present a comparison of the selected ID and OOD datasets utilized by each OOD detector. Our focus lies on experiments employing the most extensive available datasets. Additionally, we enrich this analysis by including adversarial robust OOD detectors discussed in Section 2.3. Notably, recent post-hoc detectors predominantly utilize ImageNet-1K as their standard ID datasets, whereas adversarial robust OOD detectors are more commonly evaluated on the smaller CIFAR-10 dataset. Furthermore, the architecture of adversarial robust methods diverges from lightweight post-hoc designs, often relying on adversarial training and auxiliary data, as elaborated in Section 2.3. The popular adversarial ODD are the FGSM and PGD attacks. Table 2 gives an overview of several detectors as well as considered ID and OOD datasets and model architectures. The ImageNet-1K dataset together with ResNet-50 architecture has become standard for ID. Popular OD datasets are iNaturalist, SUN, Places, and Textures. Some OOD detectors, such as ALOE, OSAD, ADT, and ATOM (see Section 2.3), aim to be robust adversarial. They usually take evading attacks (see Section 2.1) such as FGSM or PGD as OOD. These computationally expensive methods only show empirical results on the small-scale CIFAR-10. A robust OOD detector is built to distinguish whether a perturbed input is OOD. Standardized OOD benchmark frameworks, i.e. OpenOOD [5] or RoboDepth [6] do not currently include unified robustness in their benchmarks. Consequently, both frameworks, give a false sense of encompassing open-world capabilities. They focus on natural distribution [14], where OOD detection in large-scale semantic space has attracted increasing attention [71]. Some OOD datasets have issues, where ID classes are part of the OD dataset [79]. Recently, [80] found a clean semantic shift dataset that minimizes the interference of covariate shift. The experiments show that sota OOD detectors are more sensitive to covariate shifts, and the advances in semantic shift detection are minimal.
Extension to Adversarial Robust Definition
Investigating adversarial examples could gain insights into understanding the covariate shift and towards a generalized OOD detection [4]. The difference between benign ${\bm{x}}$ and the attacked counterpart ${\bm{x}}^{\prime}$ is the different attention of the pre-trained classifier $C$ per sample. Let us define an attention map $A$ [81]: The input image is passed through a classifier $C$ to obtain a feature map ${\bm{F}}$. A possible tool could be Grad-CAM [64] (see also Section 2.4) to investigate the attention change between benign ${\bm{x}}$ and adversarial example ${\bm{x}}^{\prime}$ on ${\bm{F}}$ [60] (compare Section 4.2).
Let $\Omega({\bm{x}})$ be a set of semantic-preserving perturbations on an input ${\bm{x}}$. For $\bm{\delta}\in\Omega({\bm{x}})$, ${\bm{x}}+\bm{\delta}$ have the same distributional membership (i.e., ${\bm{x}}$ and ${\bm{x}}+\bm{\delta}$ both belong to ID or OOD) [26]. To clarify, the attention maps ${\bm{A}}$ between benign ${\bm{x}}$ and adversarial example ${\bm{x}}^{\prime}$ are not the same
${\bm{A}}_{C}({\bm{x}})\neq{\bm{A}}_{C}({\bm{x}}^{\prime})\text{.}$ (3)
So adversarial examples are not from a completely different distribution, but rather adversarial perturbed versions of ID data.
@@ 4 EXPERIMENTS
In this section, we explain our experimental setup and discuss the results of 16 post-hoc OOD detectors on different datasets and pre-trained models.
4.1 Experiment Setup
We extend the OpenOOD framework [5] to consider adversarial attacks. We attack the pre-trained classifiers on the corresponding test sets and evaluate 16 post-hoc OOD detectors.
As attack methods, we choose FGSM(-$L^{\infty}$), PGD(-$L^{\infty}$), DF(-$L^{2}$) from FoolBox [82]; and the mPGD-($L^{\infty}$). The attacked models are ResNet-18 [83], ResNet-50 [83], and Swin-T [84].
The datasets are CIFAR-10 & CIFAR-100 [22], ImageNet-1K [23] and its variant ImageNet-200 with just 200 classes.
Table 3: Setup. The attack success rate (ASR) from various attacks on different models and datasets. The FGSM attack has the lowest, while DF has the highest ASR. $\mathcal{A}_{\text{std}}$ refers to the standard accuracy of the pre-trained classifier.
@@ 4.1 Experiment Setup
Figure 1: Grad-CAM comparison between the benign and its attacked counterpart. The color red indicates a high intensity of samples on similar heatmaps. The color blue indicates a low intensity. The attacks can be compared row-wise that is sorted according to datasets and attacked DNN. The DF attack yields very similar heatmaps across all datasets and models.
4.2 Grad-CAM Similarity
This section investigates the neural networks’ attention variations to classify benign and adversarial images. Our goal is to measure if adversarial attacks produce a change in the attention of the neural network that leads to misclassification as depicted in Figure 1. A previous study showed that misclassification is correlated with a shift in the center of attention for the of image classification [67]. We want to test this attention change in our pair of benign/attacked images. To quantitative measure the aforementioned attention change we use the Grad-CAM method [63]. The Grad-CAM method allows us to map the relevant part of a given image that leads to the network classification. Usually, the last block of the model is used to obtain attention [85]. In our work, we investigate the Grad-CAM output of benign and attacked images. To quantitatively analyze the Grad-CAM difference of both samples we consider two well-known metrics: the mean square error distance (L2), and the structural similar index method (SSIM) [86]. The L2 distance allows us to evaluate the pixel-wise difference of the Grad-CAM of both images. The higher the L2 value, the more distinctive the overall pixel-wise attention of the network between the benign and attacked image. The SSIM metric considers the overall structural similarity between the Grad-CAM of both images. The higher the metric the more similar the overall structure of the attention visualized through Grad-CAM maps. The combination of both metrics for our quantitative analyses is convenient because it allows us to understand better how/if an adversarial attack changes the attention of the network in an image for its classification:
  • High L2 and low SSIM report total dissimilarity of the Grad-CAM maps of the benign and attacked image. Reporting that the Grad-CAM of the attacked image is different from the benign images.
  • High L2 and high SSIM report pixel-wise differences of the Grad-CAM, but overall structural similarity. This suggests that after the attack, the attention maps shift to a different area of the image, as noted in [67].
  • Low L2 and high SSIM report total similarity, structural and pixel-wise. Reporting that the Grad-CAM is not sensitive to the changes caused by the adversarial attack.
When the L2 value approaches zero, the SSIM invariably reaches one, eliminating the possibility of simultaneously having low L2 and low SSIM. In the Figure 1, the DF attack shows a very high Grad-CAM similarity across all attack methods. Despite that, there is still a high dissimilarity by a high number of samples.